OpenStudy (anonymous):

Find the fourth roots of 16(cos 200° + i sin 200°) (check my answer).

OpenStudy (anonymous):

z1=4thrt(16)*[cos(200/4)+i*sin(200/4)]=2[cos(50)+i*sin(50)] z2=2[cos(50+90)+i*sin(50+90)]=2[cos(140)+i*sin(140)] z3=2[cos(140+90)+i*sin(140+90)]=2[cos(230)+i*sin(230)] z4=2[cos(230+90)+i*sin(230+90)]=2[cos(320)+i*sin(320)]

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Looks good.

OpenStudy (anonymous):

Thanks :)